# In Search of Adam’s Secret Sauce

### v1

@@ 5 Why an adaptive trust region? Insights from heterogeneous quadratics
Figure 7: Counterpart of Figure 5 for the heterogeneous quadratic example. We do not observe gains adding a fixed mollifier $m_{k}/\sqrt{m_{k}^{2}+\epsilon}$. Further, placing inside or outside $\sqrt{\cdot}$ has no qualitative effect.
In Proposition 1, we showed that the key difference between Signum and Adam lies in the variance correction term $\beta\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$ in the denominator. Understanding how this term evolves is essential: it cannot be approximated by a constant. In the second row of Figure 6, we observe that the variance estimate not only varies over time, but also differs in scale across the three blocks—mimicking the parameter groupings in transformer models. This block-wise variation reinforces the idea that the variance term dynamically adapts to the local curvature and cannot be substituted by a fixed value. In Figure 7 and 5, we show a similar effect in heterogeneous quadratic and language models, respectively: replacing $\beta\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$ with a fixed constant $\epsilon$ cannot provide the same adaptive effect.
@@ Acknowledgements
We would like to thank Niccolo Ajroldi, Sam Liang, and Enea Monzio Compagnoni for their valuable feedback. Antonio Orvieto acknowledges the financial support of the Hector Foundation.
@@ Appendix A Experimental details
For pre-training Transformers on Causal Language Modeling, we build upon the nanoGPT [Karpathy, 2022] implementation, augmenting it with Rotational Positional Embedding [Su et al., 2024], RMSNorm [Zhang and Sennrich, 2019], and SwiGLU [Shazeer, 2020]. All our models have a vocabolary size of 50280 and make use of GPT-Neox tokenizer [Black et al., 2022]. We adopt an enhanced training recipe, made popular by large language models such as LLaMa [Touvron et al., 2023]. These modifications include: training in bfloat16; employing a linear learning rate warm-up for 10% of the training steps, followed by cosine annealing to $51e-$ 5. Global norm clipping is used (unless specified or ablated upon) for gradients with norms above 1 (on the raw gradient, as a first step). We have no weight tying between the embedding and the last linear layer. We always report validation perplexity on a separate subset consisting of 100M tokens. Seeds, when provided, are relative to distinct network initialization.
@@ A.1 Experiments on SlimPajama – 160M parameters model
This setup comprises a total of 582 full training runs. We always use warm-up (10%) and cosine anneal until a learning rate of $51e-$ 5. This setting is Chinchilla-optimal (20 tokens/parameter).
@@
    ×$\quad\times[0.9875,0.975,0.95,0.9]$
    ×$\quad\times[0,1e-3,1e-4].$
@@
    ×$\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    ×$\quad\times[0.996875,0.99375,0.9875,0.975,0.95,0.9,0.8,0.6]$
@@
    ×$\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    ×$\quad\times[\text{True},\text{False}]$
@@
    ×$\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    ×$\quad\times[\text{True},\text{False}]$
@@
A.1.2 Sequence Length 512, Batch size 256, 3.2 B Tokens  (24800 gradient steps)
@@
A.1.3 Sequence Length 2048, Variable batch size, 2.5 B Tokens
@@
    ×$\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    ×$\quad\times[0.996875,0.99375,0.9875,0.975,0.95,0.9,0.8,0.6]$
@@
    ×$\quad\times[0.975,0.95,0.9]$
    ×$\quad\times[0.9875,0.975,0.95,0.9,0.8]$
@@ Appendix B Complementary Experimental Results
  • §B.5 confirms on the validity of our findings when ablating on nuances of Signum and Adam such as initialization and bias correction. These findings complement §3.3.
@@ B.1 Tuning for Table 1
B.2 Effect of Shorter Sequence Length in Figure 2
@@
B.3 Batch size ablation for Figure 2
@@
B.4 Figure 2 on Fineweb (no weight decay)
Finally, we evaluate our findings – both strong performance of equal $\beta$s in Adam and substantial gap with Signum on a different dataset (Fineweb [Penedo et al., 2024]). All other experiments in this paper are performed on SlimPajama. To add an additional axis of variation compared to previously presented settings, we here remove weight decay from all methods.
@@ B.4 Figure 2 on Fineweb (no weight decay)
B.5 Effect of Bias Correction and Zero Initialization on Adam
@@ C.2 Generalization of Proposition 1 – Necessity of equal betas for variance interpretation
$v_{k+1}-m_{k+1}^{2}$ $=\beta(v_{k}-m_{k+1})+\beta(1-\beta)[m_{k}^{2}+g_{k+1}^{2}-2m_{k}g_{k+1}].$
@@
$[(1-\beta_{2})-(1-\beta_{1})^{2}]g_{k+1}^{2}-2\beta_{1}(1-\beta_{1})m_{k}g_{k+1}$
@@
Substituting the definitions of Gaussian likelihood and KL divergence (6) into (4) gives the following objective function
@@
$-\lambda(g-m)-(m_{k}-m)=0\quad\Rightarrow\quad m=\frac{\lambda g+m_{k}}{1+\lambda}.$ (12)

### v2

@@ 5 Why an adaptive trust region? Insights from heterogeneous quadratics
Figure 7: Counterpart of Figure 4 for the heterogeneous quadratic example. We do not observe gains with a fixed mollifier $m_{k}/\sqrt{m_{k}^{2}+\epsilon}$. Placing inside or outside $\sqrt{\cdot}$ has no qualitative effect after tuning.
In Proposition 1, we showed that the key difference between Signum and Adam lies in the variance correction term $\beta\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$ in the denominator. Understanding how this term evolves is essential: it cannot be approximated by a constant. In the second row of Figure 6, we observe that the variance estimate not only varies over time, but also differs in scale across the three blocks—mimicking the parameter groupings in transformer models. This block-wise variation reinforces the idea that the variance term dynamically adapts to the local curvature and cannot be substituted by a fixed value. In Figure 7 and 4, we show a similar effect in heterogeneous quadratic and language models, respectively: replacing $\beta\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$ with a fixed constant $\epsilon$ cannot provide the same adaptive effect.
@@ Acknowledgements
We would like to thank Niccolo Ajroldi, Sam Liang, Weronika Ormaniec, and Enea Monzio Compagnoni for their comments. We additionally thank the NeurIPS 2025 and ICML 2025 HiLD workshop reviewers for their valuable feedback and references. Antonio Orvieto acknowledges the financial support of the Hector Foundation, and is thankful for the compute made available by MPI-IS and the Tübingen AI ecosystem.
@@ Appendix A Experimental details
For pre-training Transformers on Causal Language Modeling, we build upon the nanoGPT [Karpathy, 2022] implementation, augmenting it with Rotational Positional Embedding [Su et al., 2024], RMSNorm [Zhang and Sennrich, 2019], and SwiGLU [Shazeer, 2020]. All our models have a vocabulary size of 50280 and make use of GPT-Neox tokenizer [Black et al., 2022]. We adopt an enhanced training recipe, made popular by large language models such as LLaMa [Touvron et al., 2023]. These modifications include: training in bfloat16; employing a linear learning rate warm-up for 10% of the training steps, followed by cosine annealing to 1e51e-5. Global norm clipping is used (unless specified or ablated upon) for gradients with norms above 1 (on the raw gradient, as a first step). We have no weight tying between the embedding and the last linear layer. We always report validation perplexity on a separate subset consisting of 100M tokens. Seeds, when provided, are relative to distinct network initialization.
@@ A.1 Experiments on SlimPajama – 160M parameters model
This setup comprises a total of 747 full training runs. We always use warm-up (10%) and cosine anneal until a learning rate of 1e51e-5. This setting is Chinchilla-optimal (20 tokens/parameter).
@@
    $\quad\times[0.9875,0.975,0.95,0.9]$
    $\quad\times[0,1e-3,1e-4].$
@@
    $\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    $\quad\times[0.996875,0.99375,0.9875,0.975,0.95,0.9,0.8,0.6]$
  • Adam without global norm clipping and with $\lambda=0.1$ (weight decay) and $\epsilon=1e-8$ (usual Pytorch setup, see also Biderman et al. [2023]).
    165 full runs (Figure 12)
    $\quad\times[0.975,0.95,0.9,0.8]$
    $\quad\times[0.9875,0.975,0.95,0.9,0.8,0.6]$
@@
    $\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    $\quad\times[\text{True},\text{False}]$
@@
    $\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    $\quad\times[\text{True},\text{False}]$
@@
A.1.2 Sequence Length 2048, Batch size 256, 6.4 B Tokens  (12400 gradient steps)
The setup here is exactly as in §A.1.1, but we train for $2\times$ the token budget. We test our core claim ($\beta_{1}=\beta_{2}$ works well), and hence we run:
  • Adam with global norm clipping to 1 and with $\lambda=0.1$ (weight decay) and $\epsilon=1e-8$.
    168 full runs (Figure 13)
    $(\eta,\beta_{1},\beta_{2})$ $\in[0.032,0.016,0.008,0.004,0.002,0.001,0.0005]$
    $\quad\times[0.9875,0.975,0.95,0.9]$
    $\quad\times[0.99375,0.9875,0.975,0.95,0.9,0.8]$
A.1.3 Sequence Length 512, Batch size 256, 3.2 B Tokens  (24800 gradient steps)
@@
A.1.4 Sequence Length 2048, Variable batch size, 2.5 B Tokens
@@
    $\quad\times[0.9875,0.975,0.95,0.9,0.8]$
    $\quad\times[0.996875,0.99375,0.9875,0.975,0.95,0.9,0.8,0.6]$
@@
    $\quad\times[0.975,0.95,0.9]$
    $\quad\times[0.9875,0.975,0.95,0.9,0.8]$
@@ Appendix B Complementary Experimental Results
  • §B.6 confirms on the validity of our findings when ablating on nuances of Signum and Adam such as initialization and bias correction. These findings complement §$3.$
@@ B.1 Tuning for Table 1
Figure 12: AdamW without global norm clipping on gradients with decoupled weight decay. Compared to Figure 2, here we do not clip gradients as a first preprocessing step. Performance is slightly worse, and results are noisier. The best setting, among the ones we tried, is $\beta_{1}=\beta_{2}=0.9$. Note, however, that for large/small $\beta_{1}$s, we observe that some specific configuration with high $\beta_{2}$ can be beneficial (while still suboptimal if $\beta_{1}=\beta_{2}$ is tuned). In practice, best performance can also be achieved in this setting by merely tuning $\beta_{1}=\beta_{2}=\beta$, resulting in drastic hyperparameter grid size reduction.
B.2 Effect of More Training Tokens in Figure 2
We run part of the experiments in Figure 2 at twice the token budget. Results are conceptually very similar, and show that, on top of $\beta_{1}=\beta_{2}$ being a performance choice for AdamW, that there exists a strong correlation between $\beta$ values (see Fig. 3).
Figure 13: AdamW, same setting as Figure 2, but trained for twice the number of tokens.
B.3 Effect of Shorter Sequence Length in Figure 2
@@
B.4 Batch size ablation for Figure 2
@@
B.5 Figure 2 on Fineweb (no weight decay)
[…]
@@ B.5 Figure 2 on Fineweb (no weight decay)
B.6 Effect of Bias Correction and Zero Initialization on Adam
@@ C.2 Generalization of Proposition 1 – Necessity of equal betas for variance interpretation
$v_{k+1}-m_{k+1}^{2}$ $=\beta(v_{k}-m_{k+1}^{2})+\beta(1-\beta)[m_{k}^{2}+g_{k+1}^{2}-2m_{k}g_{k+1}].$
@@
$(1-\beta_{2})-(1-\beta_{1})^{2}]g_{k+1}^{2}-2\beta_{1}(1-\beta_{1})m_{k}g_{k+1}$
@@
Recall that
$-\log p(g_{k+1}\mid m,\sigma^{2})$ $=\frac{1}{2}\log\sigma^{2}+\frac{1}{2\sigma^{2}}(g_{k+1}-m)^{2},$
$\mathrm{KL}\left(\mathcal{N}(m_{k},\sigma^{2}_{k})\,\|\,\mathcal{N}(m,\sigma^{2})\right)$ $=\frac{1}{2}\left[\frac{\sigma^{2}_{k}}{\sigma^{2}}+\frac{(m_{k}-m)^{2}}{\sigma^{2}}-1-\log\left(\frac{\sigma^{2}_{k}}{\sigma^{2}}\right)\right].$
Therefore
$F(m,\sigma^{2})$ $=-\log p(g_{k+1}\mid m,\sigma^{2})+\frac{1}{\lambda}\,\mathrm{KL}\left(\mathcal{N}(m_{k},\sigma^{2}_{k})\,\|\,\mathcal{N}(m,\sigma^{2})\right)$
$=\frac{1}{2}\log\sigma^{2}+\frac{1}{2\sigma^{2}}(g_{k+1}-m)^{2}+\frac{1}{2\lambda}\left[\frac{\sigma^{2}_{k}}{\sigma^{2}}+\frac{(m_{k}-m)^{2}}{\sigma^{2}}-1-\log\left(\frac{\sigma^{2}_{k}}{\sigma^{2}}\right)\right]$
Since we are not optimizing for $\sigma^{2}_{k}$, we can replace $\log\left(\frac{\sigma^{2}_{k}}{\sigma^{2}}\right)=\log(\sigma^{2})$ and drop constants, gives the following objective function
@@
$\lambda(g-m)-(m_{k}-m)=0\quad\Rightarrow\quad m=\frac{\lambda g+m_{k}}{1+\lambda}.$ (13)
@@
C.4 Performance of generalized Adam reformulation
As described in §4.2, we here consider performance of the update direction:
$d_{k}=\frac{m_{k}}{\sqrt{m_{k}^{2}+\gamma\,\texttt{EMA}_{\tau}[(m_{k-1}-g_{k})^{2}]}}$ (AdaVar)
This reduces to Adam with equal betas as soon as $\beta=\gamma=\tau$ but cannot be written as an Adam update as soon as $\beta\neq\gamma$ or $\gamma\neq\tau$ (see proof in §C.2). Further, our theory in §4 shows that $\beta=\gamma=\tau$ is the only theoretically grounded choice for a precise online variational inference interpretation, also in this setting, i.e. when considering $\sigma^{2}_{k}=\gamma\,\texttt{EMA}_{\tau}[(m_{k-1}-g_{k})^{2}]$. We wonder if this insight correlates with optimal performance.
As one can see in Figure 20, we found that setting $\beta=\tau=\gamma$ leads to near optimal performance in all settings.
Figure 20: Performance of AdaVar aligns with our theoretical insights. Setup for these experiments is exactly the same as for Figure 2.